4. The elasticity of demand in the local hardware industry is -1.5, while in the video market it is -4. Which industry has a higher markup over marginal cost (as a percentage of price)? In answering, calculate the markup for each.
Answer #1

Markup over price= - 1/e

P-MC/P= -1/e

Markup of local hardware= 1/1.5= 0.67

Markup of video market= 1/4= 0.25

Thus video market has higher markup.